Let W = {(x, y) E R² | 2x + 3y = 0} Show
that W is a subspace of R2 . Find the
dimension of W. Also show that the cosets of
W are lines 2x + 3y + c = 0, where c belongs to R
1
Expert's answer
2020-12-21T17:09:40-0500

Let "W = \{(x, y) \in \mathbb R^2\ |\ 2x + 3y = 0\}". Let us show that "W" is a subspace of "\mathbb R^2". Let "(x_1,y_1),(x_2,y_2)\in W,\ a,b\in \mathbb R." Then "2x_1+3y_1=0, \ 2x_2+3y_2=0." Taking into account that "a(x_1,y_1)+b(x_2,y_2)=(ax_1+bx_2, ay_1+by_2)" and "2(ax_1+bx_2)+3(ay_1+by_2)=2ax_1+2bx_2+3ay_1+3by_2=a(2x_1+3y_1)+b(2x_2+3y_2)=a\cdot 0+b\cdot 0=0," we conclude that "a(x_1,y_1)+b(x_2,y_2)\in W", and thus "W" is a subspace of "\mathbb R^2."


For any "(x, y)\in W," "x=-1.5y", and therefore, "(x,y)=(-1.5y,y)=y(-1.5,1)". It follows that the vector "(-1.5,1)" span "W", and consequently, the dimension of "W" is 1.


The set "v+W=\{v+w\ : w\in W\}" is called a coset of "W". Let "v=(a,b)" be arbitrary element of "\mathbb R^2". Denote by "c=-2a-3b".Then "v+W=\{(a,b)+(x,y)\ | (x,y)\in W\}= \{(a+x,b+y)\ | 2x+3y=0\}=\{(x',y')\ | 2(x'-a)+3(y'-b)=0\}=\{(x',y')\ | 2x'+3y'-2a-3b=0\}=\{(x',y')\ | 2x'+3y'+c=0\}"

Therefore, the cosets of "W" are lines "2x + 3y + c = 0" , where "c\in\mathbb R."